x-3y=9
x+3y= -3
solve equation by graphing i am so lost can you help me

Found 2 solutions by ewatrrr, Alan3354:
Answer by ewatrrr(24785)

 
Hi
solve the following system by graphing)finding were the lines intersect)
x-3y=9 OR y = (1/3)x - 3 Slope = 1/3 and y intercept is Pt(0,-3) Green
x+3y= -3 OR y = (-1/3)x - 1 Slope = -1/3 and y intercept is Pt(0,-1)Blue
Lines interscet at Pt(3,-2), this ordered pair is the solution for this system
CHECKING our Answer***
x-3y=9
x+3y= -3
2x = 6, x = 3 and y = -2 (y = (1/3)x - 3 )


Answer by Alan3354(69443)

x-3y=9
x+3y= -3
solve equation by graphing
--------------
x-3y=9 --> x = 3y + 9
x+3y= -3 --> x = -3y - 3
------------
The graphs are straight lines, so 2 points are needed
x = 3y + 9
Pick 2 values for y and find x
y = 0, x = 9 --> (9,0)
y = -3, x = 0 --> (0,-3)
Plot those 2 points and draw a line thru them.
----------------
Do the same for the other equation.
Where the 2 lines intersect is the x & y values that fit both equations.
That's the solution to the system.